Iron can be extracted from haematite, Fe2O3, using either C or CO2 as the reducing agent.

The reactions are shown below.

Fe2O3(s) + 3 C(s) → 2 Fe(s) + 3 CO(g) ΔH = +492.7 kJ /mol

Fe2O3(s) + 3 CO(g) → 2 Fe(s) + 3 CO2(g) ΔH = - 24.8 kJ/mol

Substance S° (J/ K·mol)

Fe2O3 (s) 87.4

Fe (s) 27.3

C (s) 5.7

CO (g) 197.6

CO2 (g) 213.6


Use available data given to:

1. Calculate the minimum temperature at which reduction with carbon is feasible

2. Prove by calculation that reduction with carbon monoxide is feasible at all temperatures


1
Expert's answer
2022-04-26T16:59:03-0400

"\u0394G=\u0394H-T\u0394S, \u0394S1=" (3*197.6+2*27.3)-(3*5.7+87.4)=542/9 J/ K·mol, ΔH=TΔS, Tmin=ΔH/ΔS = 907.5 K

"\u0394S2= (3*213.6+2*27.3)-(3*197.6+87.4)=15.2 J\/(mol*K);" "\u0394G=\u0394H\u2212T\u0394S, \u0394G=-248000\u22120*15.2=-248000 J\/mol;" even when the temperature is 0 K the second reaction is possible, it means that reaction is feasible at all temperatures.